OpenStudy (anonymous):

A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2, and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold. Change the equation into slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all of your work. Describe how you would graph this line using the slope-intercept method.

OpenStudy (anonymous):

2x + 3y = 1200 Slope intercept form is y = mx + b. So first we subtract 2x to both sides: 3y = -2x + 1200 Now we divide 3 to both sides, can you do that?

OpenStudy (anonymous):

=400

OpenStudy (anonymous):

Yes, so we have: y = -2/3x + 400

OpenStudy (anonymous):

for part a? cause theres 6 parts :(

OpenStudy (anonymous):

Yes, that's for Part A.

OpenStudy (anonymous):

So to graph this, first we would plot the y-intercept, which is (0, 400). Then we draw a line with a slope of -2/3.

OpenStudy (anonymous):

That's Part B, where are the other parts?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

Change the equation into slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all of your work. Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences. Write the equation in function notation. Explain what the graph of the function represents. Be sure to use complete sentences. Graph the function using one of the following two options below. On the graph, make sure to label the intercepts. You may graph your equation by hand on a piece of paper and scan your work. You may graph your equation using graphic technology that can be found in the Course Information area. Suppose in the next month, the total profit on smart phone cases and tablet cases is $1,500. The profit amounts are the same, $2 for smart phone case and $3 for the tablet case. In a paragraph of at least three sentences, explain how the graphs of the functions for the two months are the same and how they are different. Be sure to use complete sentences.

OpenStudy (anonymous):

"Write the equation in function notation. Explain what the graph of the function represents. Be sure to use complete sentences." Function notation is just replacing 'y' with 'f(x)'. y = -2/3x + 400 So can you do that?

OpenStudy (anonymous):

f(x) - 2/3x + 400?

OpenStudy (anonymous):

Yes, but it's f(x) = -2/3x + 400.

OpenStudy (anonymous):

We keep the equal sign.

OpenStudy (anonymous):

Ohhh Okay Lol was that part b?

OpenStudy (anonymous):

That was for the question: "Write the equation in function notation. Explain what the graph of the function represents. Be sure to use complete sentences."

OpenStudy (anonymous):

"Graph the function using one of the following two options below. On the graph, make sure to label the intercepts." Okay, we can use technology. Let me get the graph for you.

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

There you go.

OpenStudy (anonymous):

"Suppose in the next month, the total profit on smart phone cases and tablet cases is $1,500. The profit amounts are the same, $2 for smart phone case and $3 for the tablet case. In a paragraph of at least three sentences, explain how the graphs of the functions for the two months are the same and how they are different. Be sure to use complete sentences." We can write another function, just like the last one: 2x + 3y = 1500 Let's graph them together. As you can see, the two lines are parallel, meaning they have the same slope. The only difference is that the new function has a greater y-intercept, which is 500, and the old one has a smaller y-intercept which is 400.

OpenStudy (anonymous):

That looks like all the parts :)

OpenStudy (anonymous):

Yes It Does But Dont Go Nowhere i got something else i wanted to work on with you :)

OpenStudy (anonymous):

Okay, if it's a different question, close this and open a new one :)